Startup Night 4 - 20170905 UT

Observers: JHill, BRothberg (Tucson), SWalsh (LBTO)
Telescope Operator: SAllanson (LBTO)
AO Support: D.Miller
Telescope Support: P.Hartley (LBTO)
Software Support: CBiddick, KSummers (Tucson)
Instrument Support: ESolheid (LBTO)
Engineering Support: McMahon

Plan

Pointing and Collimation Models

AGW 1&2 Transforms

SX & DX Closed Loop Function Checkout

LUCI-1/2 On-sky tests of science-like scripts

LBC Twilight Flats at end

Log

UT 03:02 - Enclosure Open. S. Allanson putting in new pointing and collimation models.

UT 03:10 - Having trouble with new models. Bare collimation and pointing model put the right primary into a limit - requested position further than can be done.

UT 03:24 - Initial input in pointing model seems to have caused the problem. S. Allanson's current values should be the starting point for the model.

UT 03:27 - Testing collimation model

UT 03:31 - AGW-1 Guideprobe seems to refuse to go on-axis. Works fine off-axis

UT 03:52 - Clouds might be interfering somewhat with tests, another possibility - absence of MODS affecting.

UT 03:57 - stopping OSS resting shells, restarting so updated config files can be ingested

UT 04:12 - Primary mirrors back to where they should be, need to see how this affects pointing/collimation tests

UT 04:40 - Pointing model works OK in azimuth @constant elevation. Change in elevation causes problems (20 degree change causes a miss in finding star). Talking with DougS we suspect that we could be using a model term that is not supported by our kernel (it has happened before). Doug will make a model using tried and true terms that we've used for many years. This could also be caused by a sign error in one of the elevation dependent terms.

UT 05:00 - adding a few more fully collimated data points for D. Summer to analyze

UT 05:41 - finished with pointing and collimation tests - moved to Stone O field for AGW 1 and 2 Transforms. Doug M. running IDL scripts as we don't have LUCI scripts for anything other than Stone L and M fields.

====================================================================================

UT 05:59 - Attempting to run Doug M.'s IDL script - troubleshooting, also possible mirror panic.

UT 06:04 - problem communicating with commitReadout.sh on LUCI computer - the exposure is injected into the observer GUI, but "UNKNOWN" appears and no exposure is taken.

UT 06:10 - nomadic thermocouple transients affecting primaries

UT 06:11 - 3sec FeII exposures for AGW imaging

UT 06:13 - sky appears to have cleared out mostly (no discernible clouds on all-sky)

UT 06:14 - mirror panic SX - calling C. Biddick

UT 06:15 - changed AGW-1 O offset to 417912 - position measured on Friday

UT 06:26 - Left Side (SX) is toast (without the primary)

UT 06:31 - Slewing to Horizon and back to see if it "jiggles" power cable back into place

UT 06:36 - Restarted scheduler to see if this fixed comm issue with Doug M's script

UT 06:37 - Restarting Scheduler Worked

- engaging LBT Right Pirate Mode

UT 06:43 - Going back to Stone O for DX AGW-2 only transform
  • - exposure 0009 3sec Dark
  • #0010 - on-axis target star
  • #0011 - restarted script - on-axis target star
  • #0012 - restarted script again - on-axis target star
  • #0013 - offset 30" - background exposure (empty)
  • probe hit limit AGW-2
  • #0014 - restarted script - on-axis target star
  • #0015 - blank sky (ish)
  • #0016 - star didn't appear - no error on my side, but error appeared on Doug M's side - looks like guide probe is in front of LUCI-2
  • #0017 - star visible
  • #0018 - 0024
  • failed to find star - might be too close to Moon

UT 07:43 - blind in LUCI-2 - redoing Stone O
  • #0025 - not sure if star is the target - seems faint
  • #0026 - background image
  • preset failed - need to redo pointing (likely the bright star in corner of mask was the actual target in #0025)
  • #0027 - on-axis star - start of new program
  • #0028 - sky image
  • #0029 - #0042
  • #0043 - on-axis star, new position angle
  • #0044 - sky image
  • #0045 - #0071
  • Ignore #0065 -was not guiding on a real star
  • #0072 - looks like guide probe in the way
UT 09:16 - finished with Stone O field - blind blind in LUCI-2

UT 09:17 - moving to another field

UT 09:28 - guidestar #23 did not appear in acquisition camera - Doug M. running field aberration off-axis tests
  • collimate on-axis, move to off-axis measure (without new collimation) - continue this in field
  • stars not found, restarting GCS
  • may have re-centered on incorrect star initially, which offset all subsequent moves
  • checked pointing and restarted
  • script is working - doing 3 stars for a basic triangular check on field aberrations
  • going to stop now and return to it when SX and DX are both operational

UT 09:55 - restarting DX AdSec - Doug M. will attempt to close the AO loop on DX as a functional checkout.

UT 10:32 - Finished functional check of DX AdSec. Doug M. closed the loop

UT 10:33 - Doing a LUCI-2 imaging test - cycle through some filters and test dither to off (guiding should stop)

UT 10:35 - Acquisition failed - no guidestar found

UT 10:37 - pointing check needed

UT 10:40 - resending preset - NGC 828 - large object to test dither to "off" field several arcminutes away - seeing ~1"
  • 5 dither H-band
  • 4 dither K-band
  • 5 dither Br-Gamma (redshifted to H2 @2.12micron for this object)
  • 5 dither H2 (redshifted to continuum @2.08micron for this object)
  • 4 dither J
  • Files: #0073-0113

UT 11:13 - seeing <1"

UT 11:40 - completed LUCI-2 imaging testing - reconfiguring for LBCs for twilight flats (clear out now)

UT 11:46 - Authorized DX only (LBC-Red) since SX Mirror is out of commission - only take flats for Red

UT 12:07 - can't figure out how to do dithering skyflats with only LBC-Red - authorizing LBC-Red only still caused fail preset

- going to zenith and taking stationary flats

UT 12:23 - Uspec and V-Bessel Flats PA=0 Uspec = 4k-8k, V = 8k-14k

UT 12:27 - Uspec and V-Bessel Flats PA=180 - last V might be ~20k

UT 12:32 - g-sloan and i-sloan PA=0 - stopping, all saturated

UT 12:35 - closing enclosure - will take LUCI-2 imaging flats

UT 12:44 - Taking LUCI-2 Imaging Flats
  • J-band - lights-ON individual read is 9k counts Halo 3 - script has 2x6.0sec x 5 exposures #0114-0123
  • H-band - lights-ON individual read is 9k count Halo 3 - script has 2x3.8sec x 5 exposures #0124-0133
  • K-band - lights ON individual read is 28k count Halo 3 (lights off is 22k) - script has 3x2.5sec x 5 exposures #0134-0143
  • BrGamma - lights ON individual read is 13k counts Halo 2 (lights off 800k) - scripts has 2x2.5 x 5 exposures #0144-0153
  • H2 - lights ON individual read is 11-12k counts Halo 2 - script has 2x2.5sec x 5 exposures #0154-0163
  • FeII - lights ON individual read is 10-11k counts Halo 2 - script has 2x2.5sec x 5 exposures #0164-0173

Summary

  • Worked on collimation and pointing - some additional insights gained - including using large offsets to blank sky with LUCI-2 later in the night - work remains (Doug S.)
  • SX Primary Mirror Acuators failed - offline for most of the night
  • AGW-2 Transforms taken by Doug M. (analysis underway)
  • Preliminary Field Aberration Data taken for DX by Doug M. (analysis underway)
  • Closed loop test on DX AO system completed (Functional Check only)
  • LUCI-2 Imaging tests done in 3 broad band filters and 2 narrow filters - tests included large offsets to blank fields - provided useful secondary test of pointing model
  • LBC Twilgiht flats done in Uspec and V-Bessell only (LBC-B and -R, respectively)
  • Internal LUCI-2 imaging flats checked for J,H,K,FeiI,Br-Gamma, and H2

-- %USERSIG{BarryRothberg - 2017-09-05}%
Topic revision: r4 - 09 Sep 2017, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback